The product of two consecutive integers is 47 more than the next consecutive integer. Find the two integers.

Answer by sextoned(4)

x(x+1) = 47+(x+2)
x^2 + x = 49 +x
subtract x from both sides leaves us with
x^2 = 49 therefore x=7 and x+1 = 8
7*8 =56
56-47=9